OpenStudy (anonymous):

what are the common factors of 12 and 15

OpenStudy (cwrw238):

12 = 2*2*3 15 = 3*5 common factor = 3

OpenStudy (anonymous):

write down the factors of both, and see what numbers the two have in common. I'll start with the easiest 12 1 * 12 15 1 * 15

OpenStudy (anonymous):

and keep going down.

OpenStudy (cwrw238):

oh yes - i forgot 1

OpenStudy (anonymous):

common factors are 1and 3